OpenStudy (anonymous):

Lane has a gift card worth $300 for a local entertainment store. Movies cost $20 each and newly released video games cost $50 each. In order to use the gift card, Lane must purchase at least seven items. Which is a combination of movies and newly released video games that Lane could purchase using his gift card? 1 movie and 6 video games 2 movies and 4 video games 6 movies and 5 video games 8 movies and 1 video game

OpenStudy (mathstudent55):

Go through each choice and add up the prices. I did the first one. Since it's more than $300, choice A is not good 1 movie and 6 video games 1 * $20 + 6 * $50 = $320 (too expensive) 2 movies and 4 video games 2 * $20 + 4 * $50 = 6 movies and 5 video games 6 * $20 + 5 * $50 = 8 movies and 1 video game 8 * $20 + 1 * $50 =

OpenStudy (anonymous):

Thanks! That makes it so much simpler!